# VIRAL: Visual Sim-to-Real at Scale for Humanoid Loco-Manipulation

### v1

@@ 2 Key Elements of VIRAL
Framework Overview To achieve efficient visual simulation training, the VIRAL controller is trained via teacher-student privileged learning [chen2020learning] as shown in Figure 2. We first train a privileged RL teacher policy with full access to the privileged state-based inputs and run the simulation without the compute burden of visual rendering on two 8-GPU L40S nodes (16 GPUs in total). During this stage, we carefully design stage-based rewards and initialize environments from demonstrations to boost RL training. Instead of training low-level motor skills from scratch, we integrate the pre-trained WBC policy [ben2025homie] and make the command for WBC the action space for the teacher policy. Details of teacher training are provided in Section 2.1.
@@ 2.1 Key Elements of Teacher Training
Teacher Formulation We formulate the teacher as a goal-conditioned RL policy. At time step $t$, the teacher $\pi_{\text{teacher}}(a_{t}|o_{t}^{\text{priv}})$ outputs a high-level command for the low-level WBC policy given privileged observation. Specifically, the teacher policy outputs $a_{t}=(\Delta\mathbf{v}_{t},\,\Delta\omega_{t}^{\text{yaw}},\Delta q^{\text{arm}}_{t},\,\Delta q^{\text{finger}}_{t})$ as the command for the WBC policy [ben2025homie], where $\Delta\mathbf{v}_{t},\,\Delta\omega_{t}^{\text{yaw}}$ are delta linear (x, y) and angular (yaw) velocity commands and $\Delta q^{\text{arm}}_{t},\,\Delta q^{\text{finger}}_{t}$ are delta joint targets for arm and finger motors. These commands are passed to the WBC policy [ben2025homie]. The privileged observation $o_{t}^{\text{priv}}=[o_{t}^{\text{prop-priv}},o_{t}^{\text{exte-priv}}]$ includes privileged proprioception and exteroception. Proprioception consists of $o_{t}^{\text{prop-priv}}=[\mathbf{v}_{t},\boldsymbol{\omega}_{t},\mathbf{g}_{t},a_{t-1},q_{t},\dot{q}_{t},\mathbf{f}^{\text{finger}}_{t}]$ where $\mathbf{v}_{t},\boldsymbol{\omega}_{t}$ are base linear and angular velocities, $\mathbf{g}_{t}$ is base projected gravity, $a_{t-1}$ is last action, $q_{t},\dot{q}_{t}$ are joint positions and velocities, $\mathbf{f}^{\text{finger}}_{t}$ are fingertip forces. As for privileged exteroception, we have $o_{t}^{\text{exte-priv}}=[e_{t},T_{t},O_{t}]$ where $e_{t}$ is the current stage, $T_{t}$ is the placement and lift target, $O_{t}$ is the relative transforms of objects and tables to the robot. All observation terms are specified in Section 7.1. The teacher is trained with PPO [schulman2017proximal] with a custom implementation of TRL [vonwerra2022trl] to train across GPUs in a distributed manner.
Teacher Element #1: Reward Design To design rewards for humanoid loco-manipulation, we segment the task into a sequence of walking, placing, grasping, and turning. Therefore, we define four key rewards:
@@
Teacher Element #2: Delta Action Space Rather than outputting absolute joint targets—as is common in legged RL locomotion [rudin2022learning]—we adopt a delta action space. The policy outputs increments that are accumulated into the WBC command. In practice, this delta representation significantly accelerates and stabilizes RL training. An ablation of this choice is shown in Figure 9.
@@
Teacher Element #3: WBC command as API To reduce reward engineering burden and enable reliable real-world deployment, the teacher in VIRAL produces high-level WBC commands rather than learning low-level motor skills from scratch. We use HOMIE [ben2025homie] as the underlying WBC controller, which provides stable lower-body locomotion and diverse upper-body poses. The command space of HOMIE involves velocity and height tracking commands for locomotion and upper-body joint commands. We extend this command interface by incorporating finger actions, yielding the full action space for VIRAL. Note that VIRAL framework does not have designs overfitting to specific WBC policy, and can be extended to other humanoid WBC controllers [ze2025twist2, luo2025sonic]. With a stable and robust WBC policy as an API layer, the action space of VIRAL policy is limited to a safe and reliable region of humanoid motions, improving deployability.
@@
Teacher Element #4: Reference State Initialization Learning long-horizon walking-placing-grasping-turning skills for high-DoF humanoids with RL typically demands heavy reward engineering still often yields suboptimal or poor sim-to-real transfer. To mitigate this, we collect 200 teleoperated simulation demonstrations and use them as a state-initialization buffer for RL (Figure 4). At every episode reset, we sample a demonstration snapshot and initialize the scene—robot, objects, and tables—accordingly, exposing the policy to diverse rewarding states long before it is capable of reaching them from scratch. This reference-biased exploration greatly reduces reliance on brittle reward tuning and improves sim-to-real transfer, as human-provided grasping and placement poses offer strong priors. Similar ideas have appeared in humanoid control [peng2018deepmimic, sharon2005synthesis] and manipulation [lin2025sim, nair2018overcoming]. We find this reset strategy to be essential for training humanoid loco-manipulation, as shown in the ablation in Figure 9.
@@ 2.2 Key Elements of Student Training
where $\rho^{o}_{\pi_{\text{teacher}}}$ and $\rho^{o}_{\pi_{\text{student}}}$ denote the observation distributions induced by the teacher and student rollouts, respectively. The distinction between DAgger and BC lies solely in the source of observations: teacher rollouts provide clean, near-optimal demonstrations that rapidly imprint strong priors on the student, while student rollouts expose the learner to states outside the teacher’s ideal distribution, which is critical for improving error-correction robustness and preventing compounding error during deployment. This mixed-policy rollout combines the fast initialization of BC with the state-coverage benefits of DAgger, producing a more resilient vision-based controller. Ablation of the mixture coefficient $\alpha$ is provided in Figure 11.
@@ Acknowledgement
We thank Jeremy Chimienti, Tri Cao, Jazmin Sanchez, Isabel Zuluaga, Jesse Yang, Caleb Geballe, Haotian Lin, Lingyun Xu, Chaitanya Chawla, Jason Liu, Tony Tao, Ritvik Singh, Ankur Handa, Arthur Allshire, Guanzhi Wang, Yinzhen Xu, Runyu Ding, Xiaowei Jiang, Yuqi Xie, Jimmy Wu, Kaushil Kundalia, Qi Wang, Scott Reed, Ziang Cao, Fengyuan Hu, Sirui Chen, Chenran Li, and Tingwu Wang for their help and support during this project.

### v2

@@ 2 Key Elements of VIRAL
Framework Overview. To achieve efficient visual simulation training, the VIRAL controller is trained via teacher-student privileged learning [chen2020learning] as shown in Figure 2. We first train a privileged RL teacher policy with full access to the privileged state-based inputs and run the simulation without the compute burden of visual rendering on two 8-GPU L40S nodes (16 GPUs in total). During this stage, we carefully design stage-based rewards and initialize environments from demonstrations to boost RL training. Instead of training low-level motor skills from scratch, we integrate the pre-trained WBC policy [ben2025homie] and make the command for WBC the action space for the teacher policy. Details of teacher training are provided in Section 2.1.
@@ 2.1 Key Elements of Teacher Training
Teacher Formulation. We formulate the teacher as a goal-conditioned RL policy. At time step $t$, the teacher $\pi_{\text{teacher}}(a_{t}|o_{t}^{\text{priv}})$ outputs a high-level command for the low-level WBC policy given privileged observation. Specifically, the teacher policy outputs $a_{t}=(\Delta\mathbf{v}_{t},\,\Delta\boldsymbol{\omega}_{t}^{\text{yaw}},\Delta\mathbf{q}^{\text{arm}}_{t},\,\Delta\mathbf{q}^{\text{finger}}_{t})$ as the command for the WBC policy [ben2025homie], where $\Delta\mathbf{v}_{t},\,\Delta\boldsymbol{\omega}_{t}^{\text{yaw}}$ are delta linear (x, y) and angular (yaw) velocity commands and $\Delta\boldsymbol{q}^{\text{arm}}_{t},\,\Delta\boldsymbol{q}^{\text{finger}}_{t}$ are delta joint targets for arm and finger motors. These commands are passed to the WBC policy [ben2025homie]. The privileged observation $o_{t}^{\text{priv}}=[o_{t}^{\text{prop-priv}},o_{t}^{\text{exte-priv}}]$ includes privileged proprioception and exteroception. Proprioception consists of $o_{t}^{\text{prop-priv}}=[\mathbf{v}_{t},\boldsymbol{\omega}_{t},\mathbf{g}_{t},\boldsymbol{a}_{t-1},\boldsymbol{q}_{t},\boldsymbol{\dot{q}}_{t},\mathbf{f}^{\text{finger}}_{t}]$ where $\mathbf{v}_{t},\boldsymbol{\omega}_{t}$ are base linear and angular velocities, $\mathbf{g}_{t}$ is base projected gravity, $\boldsymbol{a}_{t-1}$ is last action, $\boldsymbol{q}_{t},\boldsymbol{\dot{q}}_{t}$ are joint positions and velocities, $\mathbf{f}^{\text{finger}}_{t}$ are fingertip forces. As for privileged exteroception, we have $o_{t}^{\text{exte-priv}}=[e_{t},\boldsymbol{T}_{t},\boldsymbol{O}_{t}]$ where $e_{t}$ is the current stage, $\boldsymbol{T}_{t}$ is the placement and lift target, $\boldsymbol{O}_{t}$ is the relative transforms of objects and tables to the robot. All observation terms are specified in Section 7.1. The teacher is trained with PPO [schulman2017proximal] with a custom implementation of TRL [vonwerra2022trl] to train across GPUs in a distributed manner.
Teacher Element #1: Reward Design. To design rewards for humanoid loco-manipulation, we segment the task into a sequence of walking, placing, grasping, and turning. Therefore, we define four key rewards:
@@
Teacher Element #2: Delta Action Space. Rather than outputting absolute joint targets—as is common in legged RL locomotion [rudin2022learning]—we adopt a delta action space. The policy outputs increments that are accumulated into the WBC command. In practice, this delta representation significantly accelerates and stabilizes RL training. An ablation of this choice is shown in Figure 9.
@@
Teacher Element #3: WBC command as API. To reduce reward engineering burden and enable reliable real-world deployment, the teacher in VIRAL produces high-level WBC commands rather than learning low-level motor skills from scratch. We use HOMIE [ben2025homie] as the underlying WBC controller, which provides stable lower-body locomotion and diverse upper-body poses. The command space of HOMIE involves velocity and height tracking commands for locomotion and upper-body joint commands. We extend this command interface by incorporating finger actions, yielding the full action space for VIRAL. Note that VIRAL framework does not have designs overfitting to specific WBC policy, and can be extended to other humanoid WBC controllers [ze2025twist2, luo2025sonic]. With a stable and robust WBC policy as an API layer, the action space of VIRAL policy is limited to a safe and reliable region of humanoid motions, improving deployability.
@@
Teacher Element #4: Reference State Initialization. Learning long-horizon walking-placing-grasping-turning skills for high-DoF humanoids with RL typically demands heavy reward engineering still often yields suboptimal or poor sim-to-real transfer. To mitigate this, we collect 200 teleoperated simulation demonstrations and use them as a state-initialization buffer for RL (Figure 4). At every episode reset, we sample a demonstration snapshot and initialize the scene—robot, objects, and tables—accordingly, exposing the policy to diverse rewarding states long before it is capable of reaching them from scratch. This reference-biased exploration greatly reduces reliance on brittle reward tuning and improves sim-to-real transfer, as human-provided grasping and placement poses offer strong priors. Similar ideas have appeared in humanoid control [peng2018deepmimic, sharon2005synthesis] and manipulation [lin2025sim, nair2018overcoming]. We find this reset strategy to be essential for training humanoid loco-manipulation, as shown in the ablation in Figure 9.
@@ 2.2 Key Elements of Student Training
where $\rho^{o}_{\pi_{\text{teacher}}}$ and $\rho^{o}_{\pi_{\text{student}}}$ denote the observation distributions induced by the teacher and student rollouts, and $\alpha$ denotes the ratio. The distinction between DAgger and BC lies solely in the source of observations: teacher rollouts provide clean, near-optimal demonstrations that rapidly imprint strong priors on the student, while student rollouts expose the learner to states outside the teacher’s ideal distribution, which is critical for improving error-correction robustness and preventing compounding error during deployment. This mixed-policy rollout combines the fast initialization of BC with the state-coverage benefits of DAgger, producing a more resilient vision-based controller. Ablation of the mixture coefficient $\alpha$ is provided in Figure 11.
@@ Acknowledgement
We thank Jeremy Chimienti, Tri Cao, Jazmin Sanchez, Isabel Zuluaga, Jesse Yang, Caleb Geballe, Haotian Lin, Lingyun Xu, Huanyu Li, Chaoyi Pan, Chaitanya Chawla, Jason Liu, Tony Tao, Ritvik Singh, Ankur Handa, Arthur Allshire, Guanzhi Wang, Yinzhen Xu, Runyu Ding, Xiaowei Jiang, Yuqi Xie, Jimmy Wu, Haoyu Xiong, Avnish Narayan, Kaushil Kundalia, Qi Wang, Scott Reed, Ziang Cao, Fengyuan Hu, Sirui Chen, Chenran Li, and Tingwu Wang for their help and support during this project.